A Carnot engine operating at 35% efficiency exhausts heat into the heat sink at 320 ℃. If we want to increase the efficiency of the engine to 49%, without changing the temperature of the heat reservoir, what would be the new temperature of the heat sink?
Answer #1

Now, n2 =49% sol. no = 35 4. Tsink = Tz, Tsource = Ti Here Tg = 320°C 593K Using, n = 1- Tz 0.35 = 1 - 593 T :: Ti = 912.31 K